OpenStudy (anonymous):

I was born December 31st 1996 and my guy friend was born in april after me. He is 15 and I am 16. Am I only 4 months older than him??

OpenStudy (jdoe0001):

ahemm.... well... depends on when in April he was born I guess

OpenStudy (anonymous):

the 18th

OpenStudy (jdoe0001):

well, then, just 3 1/2 months you were born in december 31st, the end of the month really, just as good as 1st of january

OpenStudy (anonymous):

Well, lets see: Lets say relative to January, you were born on the 12th month; let your birthday be represented by \(b_y\). Lets write this as: \[b_y=\frac{12}{12}\] Your..."guyfriend" (haha ;)) is born in april of next year (the fourth month). Let his birthday be represented by \(b_g\). So: \[b_g=1\frac{4}{12}=\frac{16}{12}\] Your age differences, assuming you both were born on the same day is: \[b_g-b_y=\frac{16}{12}-\frac{12}{12}=\frac{4}{12}\] So therefore, assuming you were both were born on the same day, you ARE four months apart

OpenStudy (anonymous):

oh okay!! That makes sense! Thanks so much for your help

OpenStudy (anonymous):

I never thought about it until now and with my birthday being the 31st, it was confusing for me because if I was born January 1st, I would only be 15 too.

OpenStudy (anonymous):

Removing the assumption would alter the equation to be: \[b_g-b_y=4-\frac{31-d}{31}\] Where d is the date he is born. So if he is on the 18th, it would be: \[4-\frac{31-18}{31}=4-\frac{13}{31}=\frac{111}{31}=3\frac{8}{31}\] This represents three months and eight days out of A month haha that's math for ya.

OpenStudy (anonymous):

oh okay lol thanks so much. Math confuses me lol

OpenStudy (anonymous):

Hey my sister says I would actually be a year and a little over 3 months older than him. He is turning 16 in april and I am turning 17 December 31st.